# A Guide to Stochastic Optimisation for Large-Scale Inverse Problems

### v1

@@ 5.3   Fan-Beam CT
Figure 4: Reference solutions for high, medium and low noise levels
@@
$g(x)=\alpha_{I}\|\nabla x\|_{1}+\iota_{[0,\infty)^{d}}(x)$, $A=K$ with $f(x)=\frac{1}{2}\|x-v\|^{2}$ for PDHG and $A_{i}=K_{i}$ with $f_{i}(x)=\frac{1}{2}\|x-v_{i}\|^{2}$ for SPDHG, $h\equiv 0$.
The proximal operator of $g$ is approximated by 100 FGP iterations [BT09]. Stochastic methods use $n=60$ (for SGD, SAGA and SVRG), and $\ell=60$ (for SPDHG), subsets of the data. Operators $K_{i}$ are defined analogously to the parallel CT case in Section 5.2: starting with the angle $i/n$ (or $i/\ell$) we take every $n$ (or $\ell$) angle from the equidistant partition of $[0,2\pi)$ into $360$ angles. As a result, the norms $\|K_{i}\|$ are all nearly the same. SGD uses a constant step-size $\tau=1/(2nL_{\max})$, and SAGA and SVRG use $\tau=1/(3nL_{\max})$. SPDHG is implemented following [EMS19, Algorithm 2], with $\sigma=\gamma\rho/K_{\text{max}}$, $\tau=1/(n\gamma K_{\text{max}})$, with $\rho=0.99$, $\gamma=1$ and $K_{\text{max}}=\max_{i}\|K_{i}\|$.
@@
In spite of the superior convergence behaviour of SAGA, SVRG and SPDHG, running a large number of iterations might not be desirable due to computational or time constraints or even needed if the sought features of the image are recovered early. It is therefore of interest to investigate which algorithms recover a good quality reconstruction within a limited iteration budget, which image details are missing, and whether an increased computational or memory budget is justified or whether SGD would suffice. To investigate this, we look at reconstructions with SGD, SVRG, SAGA, and SPDHG within the first $10$ data passes.
@@
The results in Figure 6 confirm that all the stochastic algorithms converge and recover a visually satisfying solution quickly. Moreover, the variance reduced algorithms shows a steady convergence behaviour and better quality of detail. However, SGD also recovers a good quality solution, and depending on the problem, only a handful of data passes might be needed to recover general features of the solution. Thus, a sound strategy to improve the performance of an image reconstruction method would be to warm start the reconstruction with a few data passes with standard SGD, before employing a more advanced method. This would also help reduce the initially high variance seen with SAGA, which updates a gradient stored in the table only once the corresponding batch (or subset) index is selected (which introduces an initial bias towards selected subset gradients) and may result in a reconstruction of unreliable quality within the first $10$ data passes.
@@
As argued in Section 4, the step-size regimes are often pessimistic, in the sense that they are intended to serve as the largest value that ensures convergence towards a minimum, with a given method, for all admissible images. In Figure 7 we investigate this behaviour by looking at the convergence trajectory of SVRG for medium noise ($I=250$), with $n\in\{30,60,120\}$ and constant step-sizes $\tau=\delta/(nL_{\max})$ for $\delta\in\{1/3,1,2\}$. The last step-size was chosen such that it is the largest for which SVRG still converged to a good solution. All three batch sizes stopped converging for $\delta=7/3$. The results show that SVRG is stable with respect to the step-size choice. It gets progressively more unstable as the number of subsets $n$ increases: when $n=120$, SVRG performs best for $\delta=1$ and the worst for $\delta=2$, whereas for $n=30$ it performs the best for $\delta=2$ and the worst for $\delta=1/3$. This behaviour is (roughly) consistent across the studied noise levels. Note that the same experiment was carried out for SAGA, which showed divergent behaviour already for $\delta=1$.
@@
Figure 6: Reconstruction with SGD, SVRG, SAGA, and SPDHG for medium noise. The colour scheme in all the images is such that the smallest and largest value corresponds to the maximum and minimum values in the reference solution.
Figure 7: Behaviour of SVRG with for the reconstruction of the walnut with respect to the step-size and number of subsets $n$.
@@ 6   Outlook
Optimisation algorithms are a backbone for solving inverse problems with stochastic approaches becoming increasingly relevant for large-scale applications. We have discussed state-of-the-art algorithms for inverse problems with a particular focus on the toolbox nature of the topic. Virtually all relevant algorithms are assembled from a fairly small range of tools from both the deterministic and stochastic optimisation toolkit. Particularly the stochastic toolkit has been developed with machine learning in mind but it turns out to be of excellent use for inverse problems, too.
@@
Finally, as we already indicated throughout the manuscript in various further reading sections, an exciting path forward is the interplay between inverse problems, optimisation and machine learning. This could be in the context of modelling like learning regularisation models or learning forward operators or their sampling, but also learning optimal algorithms to solve inverse problems. We believe that the current optimisation-based paradigm with learned explicit regularisation will become a dominating computational paradigm in inverse problems, while developing efficient optimisation algorithms tailored for these applications will continue to be important to practitioners. To the best of our knowledge this has not been attempted in the context of stochastic algorithms. In our view this development is far from over and more research is needed to fully understand and exploit the connections in the context of inverse problems.
We acknowledges support from the EPSRC: MJE (EP/Y037286/1, EP/S026045/1, EP/T026693/1, EP/V026259/1) and ZK (EP/X010740/1). JL is supported by the “Fundamental Research Funds for the Central Universities”, the National Science Foundation of China (BC4190065) and the Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102).

### v2

@@ 5.3   Fan-Beam CT Data of a Walnut
Figure 4: Reference solutions for the walnut with high, medium and low noise levels
@@
$g(x)=\alpha_{I}\|\nabla x\|_{1}+\iota_{[0,\infty)^{d}}(x)$, $h\equiv 0$, $A=K$ with $f(y)=\frac{1}{2}\|y-v\|^{2}$ for PDHG and $A_{i}=K_{i}$ with $f_{i}(y)=\frac{1}{2}\|y-v_{i}\|^{2}$ for SPDHG.
The proximal operator of $g$ is approximated by 100 FGP iterations [BT09]. Stochastic methods use $n=60$ (for SGD, SAGA and SVRG), and $\ell=60$ (for SPDHG), subsets of the data. Operators $K_{i}$ are defined analogously to the parallel-beam CT case in Section 5.2: starting with the angle $i/n$ (or $i/\ell$) we take every $n$-th (or $\ell$-th) angle from the equidistant partition of $[0,2\pi)$ into $360$ angles. As a result, the norms $\|K_{i}\|$ are all nearly the same. SGD uses a constant step-size $\tau=1/(2nL_{\max})$, and SAGA and SVRG use $\tau=1/(3nL_{\max})$. SPDHG is implemented following [EMS19, Algorithm 2], with $\sigma=\gamma\rho/K_{\text{max}}$, $\tau=1/(n\gamma K_{\text{max}})$, with $\rho=0.99$, $\gamma=1$ and $K_{\text{max}}=\max_{i}\|K_{i}\|$.
@@
In spite of the superior convergence behaviour of SAGA, SVRG and SPDHG, running a large number of iterations might not be desirable due to computational or time constraints or even needed if the sought features of the image are recovered early. It is therefore of interest to investigate which algorithms recover a good quality reconstruction within a limited iteration budget, which image details are missing, and whether an increased computational or memory budget is justified or whether SGD would suffice. To investigate this, we look at reconstructions with SGD, SVRG, SAGA, and SPDHG within the first $10$ data passes for medium noise
@@
The results in Figure 6 confirm that all the stochastic algorithms perform well and recover a visually satisfying solution quickly. Moreover, variance reduced algorithms show a steady convergence behaviour and better quality of detail. SGD performs well but fails to reconstruct the desired image structures promoted by the TV regulariser and retains more reconstruction noise, within this time frame. However, depending on the problem, only a handful of data passes might be needed to recover general features of the solution. Thus, a sound strategy to improve the performance of an image reconstruction method would be to warm start the reconstruction with a few data passes with standard SGD, before employing a more advanced method. This would also help reduce the initially high variance seen with SAGA, which updates a gradient stored in the table only once the corresponding batch (or subset) index is selected (which introduces an initial bias towards selected subset gradients) and may result in a reconstruction of unreliable quality within the first $10$ data passes.
@@
As argued in Section 4, theoretical step-size regimes are often pessimistic, in the sense that they are intended to give the largest value that ensures convergence towards a minimum, with a given method, for all functions and data in the given function class. In Figure 7 we take a closer look at this behaviour by looking at the convergence trajectory of SVRG for medium noise ($I_{0}=250$), with $n\in\{30,60,120\}$ and constant step-sizes $\tau=\delta/(nL_{\max})$ for $\delta\in\{1/3,1,2\}$. SVRG with step-size $\delta=7/3$ no longer showed a convergent behaviour within the studied optimisation window, for all the investigated batch sizes. The results show that SVRG is stable with respect to the step-size choice. The variance progressively increases as the number of subsets $n$ increases: when $n=120$, SVRG performs best for $\delta=1$ and the worst for $\delta=2$, whereas for $n=30$ it performs the best for $\delta=2$ and the worst for $\delta=1/3$. This behaviour is (roughly) consistent across the studied noise levels. The same experiment was carried out for SAGA, which showed divergent behaviour already for $\delta=1$.
5.4   Image Reconstruction of Simulated PET Data
The sensitivity of the convergence behaviour of an iterative method on the Lipschitz constant can have severe consequence in reconstruction tasks where the Lipschitz constant is not available, hard to estimate, or is highly varying. A prototypical case is PET, where the data fidelity term is the Poisson log-likelihood, or equivalently, the Kullback–Leibler (KL) divergence. This data fidelity term is in general not globally Lipschitz smooth. However, in case of PET it is Lipschitz smooth for any realistic PET scan which has non-zero scatter and background events. However, the corresponding Lipschitz constant is pessimistic and usually not used in practice. Stochastic (and non-stochastic) gradient methods are nevertheless used for PET, and despite the fact that their application is in principle more heuristic, stochastic methods can show fast and reliable performance [TAK+23, KTA+21, CERS18]. To provide further insight into this behaviour, we report the numerical results from [PdCLD+24] that use the stochastic framework from the CIL library. The results concern the reconstruction of a simulated PET thorax dataset with simulated Poisson noise.
In terms of the optimisation template (7) we write the objective as follows, where $r$ denotes the background and scatter events.
Gradient based methods
$g(x)=0.1\|\nabla x\|_{1}+\iota_{[0,\infty)^{d}}(x)$, $f\equiv 0$, and $h(x)=\text{KL}(v|Kx+r)$ for FISTA and $h_{i}(x)=\text{KL}(v_{i}|K_{i}x+r_{i})$ for SGD, SAGA and SVRG;
SPDHG
$g(x)=0.1\|\nabla x\|_{1}+\iota_{[0,\infty)^{d}}(x)$, $A_{i}=K_{i}$ with $f_{i}(y)=\text{KL}(v_{i}|y+r_{i})$, $h\equiv 0$.
The numerical study compares SGD, SAGA, SVRG, Loopless-SVRG (L-SVRG) and SPDHG, therein termed Prox-SGD, Prox-SAGA, Prox-SVRG and Prox-LSVRG, respectively, against FISTA, in terms of the relative objective error and the distance to the reference solution. Stochastic methods use $32$ equidistant subsets of data. Stepsize regimes were chosen via a rough grid search, aiming to optimise the performance in terms of the distance to the solution. FISTA uses stepsize $\tau=0.0455$. SGD uses $\tau=0.1$, SAGA uses $\tau=0.025$, SVRG and L-SVRG use $\tau=0.05$, and L-SVRG uses the probability parameter $p=0.0078$. SPDHG uses $\sigma=\gamma\rho/K_{\text{max}}$, $\tau=1/(n\gamma^{2}K_{\text{max}})$, with $\rho=0.99$, $\gamma=0.25$ and $K_{\text{max}}=\max_{i}\|K_{i}\|$.
The results, reported in Figure 8a, show that all the methods show good behaviour in terms of the distance to the optimal solution, with SAGA being different compared to its performance in CT. However, in terms of the objective function error we see a somewhat different picture. Namely, SVRG, L-SVRG and SPDHG perform well, but SAGA is (qualitatively) indistinguishable from SGD and fails to show variance-reducing behaviour (which can be attributed to the fact that the variance is only expected to be reduced as we get closer to the solution). This diverging optimisation behaviour is particularly interesting due to the fact that SVRG and SAGA have the same step-size conditions in many problems, but in practice often show a different behaviour.
@@ 5.4   Image Reconstruction of Simulated PET Data
Figure 6: Reconstruction of the walnut from fan-beam CT with medium noise using SGD, SVRG, SAGA, and SPDHG. The colour scheme in all the images is such that the smallest and largest value corresponds to the maximum and minimum values in the reference solution.
Figure 7: Robustness of SVRG with respect to mis-specification of the step-size for the fan-beam reconstruction of the walnut.
(a) Convergence of iterates
(b) Convergence of objective value. Function $F$ is equal to the function $\Psi$ from template (22)
Figure 8: PET reconstruction of a 2D thorax, from [PdCLD+24]. Here Prox-SGD, Prox-SAGA, Prox-SVRG and Prox-LSVRG refer to SGD, SAGA, SVRG, Loopless-SVRG.
@@ 6   Outlook
Optimisation algorithms are a backbone for solving inverse problems and stochastic approaches becoming increasingly relevant for large-scale applications. We have discussed state-of-the-art algorithms for inverse problems with a particular focus on the toolbox nature of the topic. Virtually all of the relevant algorithms are assembled from a fairly small range of tools from both the deterministic and the stochastic optimisation toolkit. Particularly, the stochastic toolkit has been developed with machine learning in mind but it turns out to be of excellent use for inverse problems, too.
@@
Finally, as we indicated throughout the manuscript in various further reading sections, an exciting path forward is the interplay between inverse problems, optimisation and machine learning. This could be in the context of modelling, such as learning regularisation models or learning forward operators or their sampling, but also in the context of learning optimal algorithms to solve inverse problems. We believe that the current optimisation-based paradigm with learned explicit regularisation will become a dominating computational paradigm in inverse problems, while developing efficient optimisation algorithms tailored for these applications will continue to be important to practitioners. To the best of our knowledge this has not been attempted in the context of stochastic algorithms. In our view, this development is far from over and more research is needed to fully understand and exploit the connections in the context of inverse problems.
We acknowledge support from the EPSRC: MJE (EP/Y037286/1, EP/S026045/1, EP/T026693/1, EP/V026259/1) and ZK (EP/X010740/1). JL is supported by the “Fundamental Research Funds for the Central Universities”, the National Science Foundation of China (BC4190065) and the Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102).